# Weakly Supervised Co-training with Swapping Assignments for Semantic Segmentation

### v1

@@ 4.2 Ablation Studies
(a) mIoU (inc.) Base. GC SA RAW CSC DT VOC COCO 55.96 37.32 63.09 (+7.13) 42.55 (+5.23) 64.41 (+8.45) 43.92 (+6.60) 68.22 (+12.26) 45.39 (+8.07) 71.66 (+15.70) 47.10 (+9.78) 75.54 (+19.58) 49.67 (+12.35) 76.19 (+20.23) 51.00 (+13.68)
(b) mIoU (dec.) CoSA GC SA RAW CSC DT VOC COCO 76.19 51.00 75.54 (-0.65) 49.67 (-1.33) 69.89 (-6.30) 45.95 (-5.05) 72.45 (-3.74) 47.83 (-3.17) 72.10 (-4.09) 49.04 (-1.96) 74.12 (-2.07) 49.67 (-1.33)
(a) Source Detach train val GT None 83.99 80.16 NO 72.28 71.38 SPL $F$ 74.19 73.36 SPL $W_{\text{fc}}$ 78.05 76.15 SPL None 78.54 76.37
(b) Method C-mIoU mIoU FPR [8] 53.09 53.34 MCT [65] 58.46 61.24 ToCo [53] 63.62 72.33 w/o CSC 62.61 67.82 w/ CSC 82.34 76.37
Table 4: Ablation Study of GC & CSC. (a): CPL performance comparison on VOC for CAMs. Detach: stop gradient in GC for feature map $F$ or FC weights $W_{\text{fc}}$. Source: guidance sources. (b): CPL performance comparison on VOC. FPR, MCT and ToCo results are based on provided code and weights. C-mIoU: mIoU for classes with coexistence issues.
@@
Figure 8: Coexistence Problems & Effect of CSC. The class activation for bird, train, airplane, and boat are presented from left to right. (best viewed under zoom)
Impact of Guided CAMs. We evaluate the impact of including guided CAMs w.r.t. CAM quality, comparing a baseline using vanilla CAMs [72] as CPL following [70, 52] with the proposed guided CAMs. As shown in Tab. 4(a), our guided CAMs notably enhance CPL quality by 6.26% and 4.99% for train and val splits. Further, we conduct experiments to ascertain the extent to which the two CAM components, feature map $F$ and classification weights $W_{\text{fc}}$, exert greater impact on guiding CAMs. As shown, disconnection of $F$ from the gradient chains results in 74.19% and 73.36%, while the detachment of $W_{\text{fc}}$ decrease the results slightly to 78.05% and 76.37%. This suggests that guiding CAMs primarily optimizes the feature maps, verifying our initial hypothesis of the inherent non-deterministic feature map contributing to the stochasticity of CAMs in Sec. 3.1.
Impact of Swapping Assignments (SA). Tab. 3(b) suggests that eliminating SA results in significant mIoU decreases, highlighting the importance of this training strategy. Further examination of the ON and AN w.r.t. SPL and CPL indicates that, in later training stages, AN consistently outperforms ON for both SPL and CPL, as shown in Fig. 8, due to AN performing a form of model ensembling similar to Polyak-Ruppert averaging [49, 54]. We observe a noticeable disparity of mIoUs between two ONs (solid orange line vs. solid blue line in Fig. 8), which may be attributed to the superior quality of CPL and SPL from the AN facilitating a more robust ON for CoSA. The momentum framework, originally introduced to mitigate noise and fluctuations of the online learning target [22, 6], is used for information exchange across CAMs and segmentation in CoSA. To the best of our knowledge, we are the first to apply and demonstrate the efficacy of this type of training scheme in WSSS.
@@
Impact of CSC. Our CSC is introduced to address the coexistence issue. We establish C-mIoU to measure the CAM quality for those coexistence-affected classes. As shown in Tab. 4(b), applying CSC sees a boost in C-mIoU and mIoU, which surpass the existing methods. Some visual examples demonstrating these enhancements are given in Fig. 8.
Impact of Dynamic Threshold. We evaluate CoSA using some predetermined thresholds, comparing them with one employing dynamic threshold on VOC val split (see Supp. Materials for results). The performance is sensitive to the threshold, but dynamic thresholding achieves 0.65% increased performance over the best manual finetuning while saving 80% of hyper-parameter searching time.
@@ 5 Conclusion
Appendix A Additional Results
A.1 Hyper-parameter Finetuning
Here, we examine the impact of hyper-parameter variation with CoSA resulting from our finetuning. The fine-tuning of each hyper-parameter is demonstrate with the remaining parameters fixed at their determined optimal values.
Fixed Threshold vs. Dynamic Threshold. In this study, we evaluate CoSA with predetermined thresholds. The results are presented in Fig. 9. As shown, the performance peaks when this threshold is set to $0.45$, with an mIoU of $75.54\%$. However, our dynamic threshold can outperform the best manual finetuning by $0.65\%$. Despite the incurred additional $10\%$ computation overhead, our threshold searching algorithm obviates time-consuming finetuning efforts, resulting in nearly 80% reduction in hyper-parameter searching time in this case and $(1-1.1n^{-1})\%$ in general where $n$ thresholds are considered. In addition, the adoption of dynamic thresholding can enhance the generalizability to novel datasets.
(a) Seg2CAM weight $\lambda_{\text{s2c}}$ $\lambda_{\text{s2c}}$ mIoU $0.2$ 73.79 $0.1$ 74.67 $0.05$ 76.19 $0.025$ 75.25 $0.0125$ 74.33
(b) CAM2Seg weight $\lambda_{\text{c2s}}$ $\lambda_{\text{c2s}}$ mIoU $0.4$ 74.67 $0.2$ 75.56 $0.1$ 76.19 $0.05$ 73.95 $0.025$ 61.55
(c) Perplexity filter $\epsilon$ $\epsilon$ mIoU $\infty$ 73.66 $2$ 75.52 $1$ 76.19 $0.5$ 74.30 $0.1$ 70.63
(d) Momentum $m$ $m$ mIoU $0.9990$ 73.79 $0.9992$ 75.40 $0.9994$ 76.19 $0.9996$ 75.58 $0.9999$ 71.42 $1.0000$ 15.99
Table 5: Hyper-parameters Finetuning Results. Parameter searching for (a) Loss weight for CAM2Seg $\lambda_{\text{c2s}}$; (b) Loss weight for Seg2CAM $\lambda_{\text{s2c}}$; (c) Low-pass perplexity filter coefficient $\epsilon$; (e) EMA Momentum $m$ for updating assignment network. mIoU represents semantic segmentation result on PASCAL VOC val split.
A.2 Per-class Segmentation Comparisons
Comparisons on VOC. Tab. 12 illustrates the CoSA per-class mIoU results compared with recent works: AdvCAM [33], MCT [65], ToCo [53], Xu et al. [66], BECO [50]. To be fair in comparison, we include CoSA with CRF [10] postprocessing results, denoted as $*{}^{*}$, same as other SOTA models. Notably, CoSA dominates in 10 out of 21 classes. In particular, categories like boat ($5.9\%\!\uparrow$), chair ($8.2\%\!\uparrow$), and sofa ($17.2\%\!\uparrow$), demonstrate substantial lead over the SOTA models. In the VOC test split (depicted in Tab. 12), we still observe its superiority over other SOTA methods, where CoSA dominates in 15 out of 21 classes.
A.3 Further Qualitative Comparisons
Appendix B Further Analysis
@@
Impact of CRF. The conditional random field (CRF) proposes to optimize the segmentation by utilizing the low-level information obtained from the local interactions of pixels and edges [10]. Traditionally, a manually designed CRF postprocessing step has been widely adopted for refining segmentation [15, 50] or CAMs [51, 65, 70] in WSSS. As our aim is to develop a fully end-to-end WSSS solution, incorporating CRF postprocessing contradicts this principal. Through our experiments, we demonstrate that CoSA, unlike other single-stage methods, does not heavily depend on CRF. Our results indicate that incorporating CRF results in marginal improvement of $0.2\%$, $0.1\%$, and $0.1\%$ for VOC val, VOC test, and COCO val, respectively, as presented in Tab. 2 of the main paper. Tab. 6(a) suggests that in comparison to other SOTA models, our CoSA exhibits a lesser dependency on the CRF postprocessing. On the contrary, eliminating the CRF step leads to a noteworthy enhancement of 165% in terms of inference speed, as demonstrated in Tab. 6(b).
(a) Method w/o CRF w/ CRF AFA [52] 63.8 66.0 (+2.2) VIT-PCM [51] 67.7 71.4 (+3.7) ToCo [53] 69.2 71.1 (+0.9) CoSA 76.2 76.4 (+0.2)
(b) CoSA Speed w/o CRF 4.11 imgs/s w/ CRF 1.83 imgs/s
@@
Table 6: CRF Impact. (a): Comparisons of CRF impact on SOTA single-stage WSSS methods on VOC val. (b): Inference speed with and without CRF. Speed tested using a single 3090 GPU.
@@
Transformation
Description
Parameter Setting
RandomRescale
Rescale the image by $r$ times, $r$ randomly sampled from $r\sim U(r_{min},r_{max})$.
$r_{min}=0.5,r_{max}=2$
RandomFlip
Randomly horizontally flip a image with probability of $p$.
$p=0.5$
RandomCrop
Randomly crop a image by a hight $h$ and a width $w$.
$w=448,h=448$
GaussianBlur
Randomly blur a image with probability of $p$.
$p=0.5$
Transformation
Description
Parameter Setting
RandomRescale
Rescale the image by $r$ times, $r$ randomly sampled from $r\sim U(r_{min},r_{max})$.
$r_{min}=0.5,r_{max}=2$
RandomFlip
Randomly horizontally flip a image with probability of $p$.
$p=0.5$
RandomCrop
Randomly crop a image by a hight $h$ and a width $w$.
$w=448,h=448$
GaussianBlur
Randomly blur a image with probability of $p$.
$p=0.5$
OneOf
Select one of the transformation in a transformation set $T$.
$T=$ TransAppearance
Transformation
Description
Parameter Setting
Identity
Returns the original image.
Autocontrast
Maximizes the image contrast by setting the darkest (lightest) pixel to black (white).
Equalize
Equalizes the image histogram.
RandSolarize
Invert all pixels above a threshold value $T$.
$T\in U(0,1)$
RandColor
Adjust the color balance. $C=0$ returns a black&white image, $C=1$ returns the original image.
$C\in U(0.05,0.95)$
RandContrast
Adjust the contrast. $C=0$ returns a solid grey image, $C=1$ returns the original image.
$C\in U(0.05,0.95)$
RandBrightness
Adjust the brightness. $C=0$ returns a black image, $C=1$ returns the original image.
$C\in U(0.05,0.95)$
RandSharpness
Adjust the sharpness. $C=0$ returns a blurred image, $C=1$ returns the original image.
$C\in U(0.05,0.95)$
RandPolarize
Reduce each pixel to $C$ bits.
$C\in U(4,8)$
Efficiency Study. Unlike multi-stage approaches, CoSA is extremely efficient in training. It can be trained end-to-end efficiently. When training a semantic segmentation model with weak labels on the VOC dataset, our method requires a mere 8.7 hours of training time and a total of 92M parameters. In contrast, MCT [65] would necessitate approximately 231% more time (20.1hrs $\uparrow$) and 173% more parameters (159M $\uparrow$) for the same task, and BECO [50] would require around 240% more time (20.9hrs $\uparrow$) and 50% more parameters (46M $\uparrow$). When compared to the single-stage method, CoSA also demonstrate its advantage in speed-accuracy trade-off. Further details regarding the efficiency study can be found in Tab. 7.

### v2

@@ 4.2 Ablation Studies
Source Detach train val GT None 83.99 80.16 NO 72.28 71.38 SPL $F$ 74.19 73.36 SPL $W_{\text{fc}}$ 78.05 76.15 SPL None 78.54 76.37
Method C-mIoU mIoU FPR [9] 53.09 53.34 ToCo [53] 63.62 72.33 SeCo [69] 73.18 73.63 w/o CSC 62.61 67.82 w/ CSC 82.34 76.37
mIoU(%) Speed Use CRF? BECO-R101 72.1 70.9(-1.2) 1.95 4.94 COSA-R101 76.5 76.4(-0.1) 2.36 9.60 ToCo-ViT 71.1 69.2(-1.9) 1.82 3.99 COSA-ViT 76.4 76.2(-0.2) 1.83 4.11
Table 4: Ablation Study of GC, CSC and CRF. (a): CPL performance comparison on VOC. Source: guidance sources. Detach: stop gradient in GC for feature map $F$ or $W_{\text{fc}}$. (b): CPL performance comparison. FPR, ToCo and SeCo results are based on their code and weights. C-mIoU: mIoU for classes with coexistence. (c): CRF Impact. Best speed-accuracy tradeoff is achieved without using CRF. Inference speeds (FPS) are tested on RTX 3090.
@@
Impact of Guided CAMs. Our model is compared with a baseline [73] that directly uses CAMs as CPL. As shown in Tab. 4(a), our guided CAMs notably enhance CPL quality by 6.26% and 4.99% for train and val splits. Further, we conduct experiments to ascertain the extent to which the two CAM components, feature $F$ and weights $W_{\text{fc}}$, exert greater impact on guiding CAMs. As shown, the deteachment of $F$ results in 74.19% and 73.36%, but $W_{\text{fc}}$ can decrease the results slightly to 78.05% and 76.37%. This suggests that guiding CAMs primarily optimizes feature maps, verifying our hypothesis of the non-deterministic feature map contributing to the stochasticity of CAMs in Sec. 3.1.
Impact of Swapping Assignments (SA). Tab. 3(b) suggests that eliminating SA results in significant mIoU decreases, highlighting the importance of this training strategy. Further examination of the ON and AN w.r.t. SPL and CPL indicates that, in later training stages, AN consistently outperforms ON for both SPL and CPL, as shown in Fig. 9, due to AN performing a form of model ensembling similar to Polyak-Ruppert averaging [49, 54]. We observe a noticeable disparity of mIoUs between two ONs (solid orange line vs. solid blue line in Fig. 9), which may be attributed to the superior quality of CPL and SPL from the AN facilitating a more robust ON for CoSA. The momentum framework, originally introduced to mitigate noise and fluctuations of the online learning target [22, 6], is used for info exchange across CAMs and segmentation in CoSA.
@@
Impact of CSC. Our CSC is introduced to address the coexistence issues. We establish C-mIoU to measure the CAM quality for those coexistence-affected classes. As shown in Tab. 4(b), applying CSC sees a boost in C-mIoU and mIoU, which surpass the existing methods. Some visual comparisons are given in Fig. 7.
Impact of Dynamic Threshold. We evaluate CoSA using some predetermined thresholds, comparing them with one employing dynamic threshold on VOC val split (see Supp. for results). The performance is sensitive to the threshold, but dynamic thresholding achieves 0.65% increased performance over the best manual finetuning while saving 80% of hyper-parameter searching time.
4.3 Further Analysis
Training and Inference Efficiency Analysis. Unlike multi-stage approaches, CoSA can be trained end-to-end efficiently. Compared to BECO [50], our method is 240% faster in training, uses 50% fewer parameters, and yields a 4.3% higher mIoU on VOC test. Please refer to Supp. for more discussion. At inference time, we find that CRF post-processing, which is commonly adopted for refining masks [15, 50] or the CAMs [51, 65, 73], can greatly slow down the inference speed. As our aim is to develop a fully end-to-end WSSS solution, incorporating CRF post-processing contradicts this principal. Through our experiments, we show that CoSA does not heavily depend on CRF: incorporating CRF results in marginal improvement of $0.2\%$, $0.1\%$, and $0.1\%$ for VOC val, VOC test, and COCO val, respectively (Tab. 1). Conversely, eliminating CRF in CoSA can greatly speed up inference (a noteworthy 307% and 165% $\uparrow$) and achieve better speed-accuracy tradeoff as suggested in Tab. 4(c).
Hyper-parameter Sensitivity. We apply grid search strategy to explore the hyper-parameters in CoSA. The analysis of parameters, such as perplexity filter, loss weights, momentum are discussed in Supp. CoSA maintains consistent with variations of those parameters, which demonstrate its robustness.
@@ 5 Conclusion
Acknowledgments
The work is supported by European Research Council under the European Union’s Horizon 2020 research and innovation programme (GA No 787768).
Appendix 0.A Further Analysis
Contrastive Separation Analysis. Fig. 10 shows the analysis of contrastive separation on COCO. A similar trend between $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ is observed as on VOC. This suggests that the distinct relationship between $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ extends beyond the VOC dataset to encompass a broader dataset as well.
Figure 10: $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ Comparisons on COCO. (left): mIoU vs. time-steps for $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ on COCO val. (right): same as (left) but filtered by perplexity.
CPL Analysis. While the proposed dynamic learning techniques in Sec. 3.3, namely reliability based adaptive weighting and dynamic threshold, were originally inspired by the VOC dataset, we demonstrate their applicability on COCO here. As shown in Fig. 11(left), the negative correlation between perplexity and accuracy remains significant across various training time-steps, highlighting our purposed perplexity estimation method can be used to penalize inaccurate regions in our CAM2Seg loss. Fig. 11(right) also suggests that the confidence distribution on the COCO dataset shows discernible clusters.
Figure 11: CPL Analysis on COCO (left) correlation between perplexity and accuracy of CPL for different time-steps. (right) distribution of CAMs’ confidence categorized by the proposed dynamic threshold
@@ Appendix 0.A Further Analysis
Impact of CRF. The conditional random field (CRF) proposes to optimize the segmentation by utilizing the low-level information obtained from the local interactions of pixels and edges [8]. Traditionally, a manually designed CRF postprocessing step has been widely adopted for refining segmentation [15, 50] or CAMs [51, 65, 73]. As our aim is to develop a fully end-to-end WSSS solution, incorporating CRF postprocessing contradicts this principal. Through our experiments, we demonstrate that CoSA, unlike other single-stage methods, does not heavily depend on CRF. Our results indicate that incorporating CRF results in marginal improvement of $0.2\%$, $0.1\%$, and $0.1\%$ for VOC val, VOC test, and COCO val, respectively, as presented in Tab. 2 of the main paper. Tab. 5(a) suggests that in comparison to other SOTA models, our CoSA exhibits a lesser dependency on the CRF postprocessing. On the contrary, eliminating the CRF step leads to a noteworthy enhancement of 165% in terms of inference speed, as demonstrated in Tab. 5(b).
Method w/o CRF w/ CRF AFA [52] 63.8 66.0 (+2.2) VIT-PCM [51] 67.7 71.4 (+3.7) ToCo [53] 69.2 71.1 (+0.9) SeCo [69] 72.2 73.7 (+1.5) CoSA 76.2 76.4 (+0.2)
CoSA Speed w/o CRF 4.11 imgs/s w/ CRF 1.83 imgs/s
@@
Table 5: CRF Impact. (a): Comparisons of CRF impact on SOTA single-stage WSSS methods on VOC val. (b): Inference speed with and without CRF. Speed tested using a RTX 3090.
Training Efficiency Study. Unlike multi-stage approaches, CoSA is extremely efficient in training. It can be trained end-to-end efficiently. When training a semantic segmentation model with weak labels on the VOC dataset, our method requires a mere 8.7 hours of training time and a total of 92M parameters. In contrast, MCT [65] would necessitate approximately 231% more time (20.1hrs $\uparrow$) and 173% more parameters (159M $\uparrow$) for the same task, and BECO [50] would require around 240% more time (20.9hrs $\uparrow$) and 50% more parameters (46M $\uparrow$). When compared to the single-stage method, CoSA also demonstrate its advantage in speed-accuracy trade-off. Further details regarding the efficiency study can be found in Tab. 6.
@@
Appendix 0.B Further Implementation Details
Appendix 0.C Additional Results
0.C.1 Contribution of each Component
In addition to the module analysis in Sec. 4.2, we further present additional ablation studies regarding the impact of each component on the baseline. A one-stage WSSS framework, proposed in RRM [73], is utilized as the baseline model, albeit without the inclusion of the offline refinement module. As illustrated in Tab. 7, it is evident that all components proposed in CoSA exhibit positive effects on the baseline model.
mIoU @ val (increase)
Base. GC SA RAW CSC DT VOC COCO
55.96 37.32
63.09 (+7.13) 42.55 (+5.23)
64.41 (+8.45) 43.92 (+6.60)
67.04 (+11.08) 42.65 (+5.33)
69.24 (+13.27) 44.11 (+6.79)
61.80 (+5.84) 43.28 (+5.96)
Table 7: Contribution of Each Component. We systematically include the proposed components on the baseline (Base.). GC: Guided CAMs, SA: Swapping Assignments, RAW: Reliability based Adaptive Weighting, CSC: Contrastive Separation in CAMs, and DT: Dynamic Threshold.
0.C.2 Hyper-parameter Finetuning
we examine the impact of hyper-parameter variation with CoSA resulting from our finetuning. The fine-tuning of each hyper-parameter is demonstrate with the remaining parameters fixed at their determined optimal values. Those hyper-parameters were tuned on VOC val set, and the SOTA results on COCO were achieved with the same hyper-parameters, except batch size and number of iterations were tweaked to accommodate the dataset size difference.
Fixed Threshold vs. Dynamic Threshold. We evaluate CoSA with predetermined thresholds. The results are presented in Fig. 12. As shown, the performance peaks when this threshold is set to $0.45$, with an mIoU of $75.54\%$. However, our dynamic threshold can outperform the best manual finetuning by $0.65\%$. Despite the incurred additional $10\%$ computation overhead, our threshold searching algorithm obviates time-consuming finetuning efforts, resulting in nearly 80% reduction in hyper-parameter searching time in this case and $(1-1.1n^{-1})\%$ in general where $n$ thresholds are considered. In addition, the adoption of dynamic thresholding can enhance the generalizability to novel datasets.
(a) Seg2CAM weight $\lambda_{\text{s2c}}$
$\lambda_{\text{s2c}}$ mIoU
$0.2$ 73.79
$0.1$ 74.67
$0.05$ 76.19
$0.025$ 75.25
$0.0125$ 74.33
(b) CAM2Seg weight $\lambda_{\text{c2s}}$
$\lambda_{\text{c2s}}$ mIoU
$0.4$ 74.67
$0.2$ 75.56
$0.1$ 76.19
$0.05$ 73.95
$0.025$ 61.55
(c) Perplexity filter $\epsilon$
$\epsilon$ mIoU
$\infty$ 73.66
$2$ 75.52
$1$ 76.19
$0.5$ 74.30
$0.1$ 70.63
(d) Momentum $m$
$m$ mIoU
$0.9990$ 73.79
$0.9992$ 75.40
$0.9994$ 76.19
$0.9996$ 75.58
$0.9999$ 71.42
$1.0000$ 15.99
Table 8: Hyper-parameter Finetuning Results. Parameter searching for (a) Loss weight for CAM2Seg $\lambda_{\text{c2s}}$; (b) Loss weight for Seg2CAM $\lambda_{\text{s2c}}$; (c) Low-pass perplexity filter coefficient $\epsilon$; (e) EMA Momentum $m$ for updating assignment network. mIoU represents semantic segmentation result on PASCAL VOC val split.
0.C.3 Per-class Segmentation Comparisons
Comparisons on VOC. Tab. 10 illustrates the CoSA per-class mIoU results compared with recent works: AdvCAM [33], MCT [65], ToCo [53], Xu et al. [66], BECO [50]. To be fair in comparison, we include CoSA with CRF [8] postprocessing results, denoted as CoSA, same as other SOTA models. Notably, CoSA dominates in 10 out of 21 classes. In particular, categories like boat ($5.9\%\!\uparrow$), chair ($8.2\%\!\uparrow$), and sofa ($17.2\%\!\uparrow$), demonstrate substantial lead over the SOTA models. In the VOC test split (depicted in Tab. 10), we still observe its superiority over other SOTA methods, where CoSA dominates in 15 out of 21 classes.
Method bkg plane bike bird boat bottle bus car cat chair cow
AdvCAM [34] CVPR21 90.0 79.8 34.1 82.6 63.3 70.5 89.4 76.0 87.3 31.4 81.3
MCT [65] CVPR22 91.9 78.3 39.5 89.9 55.9 76.7 81.8 79.0 90.7 32.6 87.1
ToCo [53] CVPR23 91.1 80.6 48.7 68.6 45.4 79.6 87.4 83.3 89.9 35.8 84.7
Xu et al. [66] CVPR23 92.4 84.7 42.2 85.5 64.1 77.4 86.6 82.2 88.7 32.7 83.8
BECO [50] CVPR23 91.1 81.8 33.6 87.0 63.2 76.1 92.3 87.9 90.9 39.0 90.2
CoSA* (Ours) 93.1 85.5 48.5 88.7 70.0 77.6 90.4 86.4 90.3 47.2 88.7
Method table dog horse mbike person plant sheep sofa train tv mIoU
AdvCAM [34] CVPR21 33.1 82.5 80.8 74.0 72.9 50.3 82.3 42.2 74.1 52.9 68.1
MCT [65] CVPR22 57.2 87.0 84.6 77.4 79.2 55.1 89.2 47.2 70.4 58.8 71.9
ToCo [53] CVPR23 60.5 83.7 83.7 76.8 83.0 56.6 87.9 43.5 60.5 63.1 71.1
Xu et al. [66] CVPR23 59.0 82.4 80.9 76.1 81.4 48.0 88.2 46.4 70.2 62.5 72.2
BECO [50] CVPR23 41.6 85.9 86.3 81.8 76.7 56.7 89.5 54.7 64.3 60.6 72.9
CoSA* (Ours) 54.1 87.3 87.1 79.6 85.6 53.2 89.9 71.9 65.1 63.4 76.4